### v3

# Phonon-driven multipolar dynamics in a spin-orbit coupled Mott insulator

Abstract
Motivated by advances in pump-probe experiments and light-driven phenomena, we theoretically study the impact of pumped and driven phonons in Mott insulators which host multipole moments, thus going beyond conventional dipolar magnetism. As a case study, we examine pseudospin-1/2 Mott insulators hosting quadrupolar and octupolar moments, and investigate the effect of resonantly exciting ${\cal E}_{g}$ phonon modes which couple linearly to the quadrupoles. We show that this leads to multipolar precession, with the back action resulting in chiral phonon dynamics below the octupolar phase transition. We further explore the impact of a coherent two-phonon Floquet drive, showing that it can be used to ‘train’ octupolar order or even switch its sign on ultrafast timescales. Our results are obtained combining a generalized Monte Carlo code incorporating phonons with molecular dynamics simulations which numerically integrate the coupled spin-phonon equations of motion. Our work shows how phonon-driven dynamics can be used to probe and control hidden orders in solids.
preprint: APS/123-QED
Introduction.— Shaking electrons or atoms with light has been proposed to induce remarkable non-equilibrium many body states in solid state crystals as well as ultracold atoms [1, 2, 3, 4, 5, 6, 7]. In recent years, rapid advances in ultrafast and THz optics have led to the ability to resonantly excite specific orbital transitions and phonon modes in solids [8, 9, 10], providing a distinct pathway to control broken symmetry states including ferroelectricity [11], superconductivity [12], and magnetism [13, 9, 14, 15].
In dipolar magnets, local Einstein phonons can strongly impact the single-ion anisotropy or spin-bilinear exchange interaction on bonds. This has been shown to stabilize unconventional equilibrium magnetic orders in geometrically frustrated lattices such as kagome, triangular, or pyrochlore systems [16, 17, 18, 19]. Going beyond equilibrium orders, recent experimental and theoretical work has shown that coherent phonon driving can produce fascinating non-equilibrium Floquet dynamics in spin-phonon coupled dipolar magnets. These include the generation of anomalously large effective magnetic fields from driven chiral phonons [20, 21, 22, 23, 24, 25, 26], tuning of magnetic orders and magnon band topology [27, 9, 14, 28], frustration breaking in dimerized magnets [29], nonequilibrium phase transitions [30], and high harmonic generation [31]. These advances have motivated the study of many-body coupled spin-phonon dynamics in such systems while incorporating dissipation effects. Earlier work has shown the value of such an approach for understanding pump-probe dynamics in layered van der Waals magnets [32].
Going beyond the physics of dipolar spin systems, there is a rich set of issues one encounters when one starts to explore multipolar degrees of freedom which are found in solids with strong spin-orbit coupling. Unlike dipole moments, higher order multipoles such as quadrupoles and octupoles are associated with spin-orbital entangled charge or current distributions which do not simply couple to conventional probes; for this reason, multipolar broken symmetries, found in diverse materials, are hard to detect and generically termed ‘hidden orders’ [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46]. Quadrupolar moments, however, being even under time-reversal symmetry, can linearly couple to phonons, which can lead to a significant differences in the equilibrium ordering and excitations of multipoles when compared with magnetic dipoles. This is exemplified by the recent understanding of unusual patterns of multipolar symmetry breaking and lattice displacements in Ba2MgReO6 [47, 48, 49, 50, 51] and the direct detection of vibronic modes by observing spectral splitting and side-bands in resonant inelastic X-ray scattering in K2IrCl6 [52].
Recent experiments have begun to extend optical and THz techniques to study the non-equilibrium dynamics of multipolar orders. For instance, resonantly excited phonon modes, on ultrafast timescales, in pump-probe and THz experiments have shown evidence of multipolar modes coupled to lattice vibrations [29]. In $\rm{Ca_{2}RuO_{4}}$, THz experiments have revealed a ‘hidden’ quadrupolar-ordered state through the anomalous broadening of coherently driven phonons [53]. Such coupling between phonons and quadrupole moments suggests that it may be possible to not only probe but also achieve the dynamical control of multipolar orders using optical and THz techniques. These developments provide an impetus for the theoretical exploration of pumped and driven multipolar materials, and call out for a theory of many-body dynamics in coupled multipolar-phonon systems.
In this Letter, we fill this important gap by developing a simulation tool to study multipolar orders and their semiclassical dynamics including the crucial effect of coupling to local Einstein phonons and incorporating dissipation effects. In order to do this, we combine large scale equilibrium spin-phonon Monte Carlo (MC) simulations with spin-phonon molecular dynamics (MD) obtained by integrating their coupled equations of motion. The key conceptual idea underlying our work is that a phonon mode coupled to a quadrupole operator acts as a transverse field to all other multipole operators which do not commute with the relevant quadrupole. Kicking the phonon thus engenders nontrivial precessional dynamics in multipolar space which can be studied theoretically and experimentally, enabling one to shed light on complex hidden orders.
While our approach is broadly applicable to many multipolar materials, we illustrate the power of this tool by focussing on a recently discovered paradigmatic example of multipolar systems in Mott insulators, the osmate double perovskites [54, 55, 56, 46, 57, 58] which have been recently shown to display competing octupolar and quadrupolar ordering tendencies [55, 56, 58]. Our work shows how the presence of local phonons, resonantly excited by either pump pulses or Floquet driving, can influence the dynamics of the multipolar moments in such a way as to enable the detection, training, and ultrafast switching of exotic octupolar order. Our results point to a number of future research directions, including experiments to probe and control elusive hidden orders using phonons.
Model.— The simplest example of multipolar moments arise in materials with non-Kramers doublets, local pseudospin-1/2 degrees of freedom whose doublet degeneracy is protected by crystalline point group symmetries rather than just time-reversal symmetry as in the case of Kramers doublets. The ability to split this doublet by breaking the local point group symmetry shows that these doublets have quadrupolar character and are not pure dipoles. Remarkably, non-Kramers doublets are found in a diverse array of quantum materials including: quantum spin-ice compounds such as Pr2Zr2O7, Pr2Sn2O7 [59, 60, 61, 62], Kondo lattice intermetallics PrTi2Al20, PrV2Al20 [63, 64, 42, 65, 43], Mott insulating double perovskites $\rm{Ba_{2}CaOsO_{6}}$, $\rm{Ba_{2}MgOsO_{6}}$, $\rm{Ba_{2}ZnOsO_{6}}$ [54, 55, 46], and vacancy ordered double perovskites Cs2WCl6 and Rb2WCl6 [66, 67].
We focus here on non-Kramers doublets with time-reversal even (${\cal T}$-even) quadrupolar moments represented by pseudospin-1/2 Pauli operators $(\tau_{x},\tau_{z})$ and a ${\cal T}$-odd octupole represented by $\tau_{y}$ as realized in the double perovskites and the Kondo intermetallics. In the double perovskite Mott insulators, these pseudospins live on a face-centered cubic (fcc) lattice, and descend from the splitting of a $J=2$ angular momentum multiplet, with $\tau_{x}\propto(J_{x}^{2}-J_{y}^{2})$, $\tau_{z}\propto(3J_{z}^{2}-J(J+1))$, and $\tau_{y}\propto\overline{J_{x}J_{y}J_{z}}$, with bar denoting operator symmetrization.
Symmetries dictate the following fcc lattice Hamiltonian describing nearest-neighbor pseudospin interactions
$\!\!\mathfrak{h}_{\rm sp}\!$ $=$ $\!\!\!\sum_{\langle{\bf{r}},{\bf{r}}^{\prime}\rangle}\!\!\!\left[-J_{0}\tau_{{\bf{r}}y}\tau_{{\bf{r}}^{\prime}y}\!+\!J_{1}\left(\cos^{2}\!\phi_{{\bf{r}}{\bf{r}}^{\prime}}\!+\!\gamma\sin^{2}\!\phi_{{\bf{r}}{\bf{r}}^{\prime}}\right)\tau_{{\bf{r}}z}\tau_{{\bf{r}}^{\prime}z}\right.$ (1)
$+$ $J_{1}\left(1-\gamma\right)\sin\phi_{{\bf{r}}{\bf{r}}^{\prime}}\cos\phi_{{\bf{r}}{\bf{r}}^{\prime}}\left(\tau_{{\bf{r}}x}\tau_{{\bf{r}}^{\prime}z}+\tau_{{\bf{r}}z}\tau_{{\bf{r}}^{\prime}x}\right)$
$+$ $\left.J_{1}\left(\sin^{2}\phi_{{\bf{r}}{\bf{r}}^{\prime}}+\gamma\cos^{2}\phi_{{\bf{r}}{\bf{r}}^{\prime}}\right)\tau_{{\bf{r}}x}\tau_{{\bf{r}}^{\prime}x}\right],$
where $\phi_{ij}=\{0,2\pi/3,4\pi/3\}$ correspond to nearest neighbors $\langle{\bf{r}}{\bf{r}}^{\prime}\rangle$ in the $\{XY,YZ,ZX\}$ planes. The phase diagram of this model depends on the dimensionless ratios of exchange couplings via $J_{1}/J_{0}$ and $\gamma$ [68, 56].
Based on density functional theory and dynamical mean field theory for Ba2CoOsO6 [69], we set $\gamma\!\approx\!-0.4$ and $J_{1}/J_{0}\!\approx\!0.5$ for the pseudospin Hamiltonian; this places us in the ferro-octupolar phase for $T/J_{0}\lesssim 3.2$ as determined from classical Monte Carlo simulations [56]. A microscopic calculation of the octupolar exchange yields $J_{0}\approx 1$ meV, which leads to $T_{\rm oct}\approx 40$ K, in good agreement with experimental results on Ba2CoOsO6 [54, 56, 46]. Motivated by constructing the more general phase diagram of this model, we fix $\gamma$ and vary $J_{1}/J_{0}$ and $T/J_{0}$ which allows us to access both ferro-octupolar and quadrupolar ordered phases.
Next, we consider cubic symmetry ${\cal E}_{g}$ Einstein phonon modes at each lattice site with $x\equiv x^{2}\!-\!y^{2}$, $z\equiv 3z^{2}\!-\!r^{2}$, with the phonon Hamiltonian
$\mathfrak{h}_{\rm{ph}}=\sum_{{\bf{r}},\alpha=x,z}\left[\frac{1}{2m}\mathfrak{p}_{\alpha}^{2}({\bf{r}})+\frac{1}{2}m\omega^{2}\mathfrak{q}_{\alpha}^{2}({\bf{r}})\right]$ (2)
where $(\mathfrak{q}_{\alpha},\mathfrak{p}_{\alpha})$ are the normal mode coordinates and momenta, and $\omega,m$ represent the frequency and mass, respectively. On symmetry grounds, the ${\cal T}$-odd octupole will not linearly couple to phonons, however the $(\tau_{x},\tau_{z})$ quadrupoles can couple linearly to the phonon modes. The symmetry allowed spin-phonon coupling is given by
$\mathfrak{h}_{\rm sp-ph}=-g\sum_{\bf{r}}\left[\mathfrak{q}_{x}({\bf{r}})\tau_{{\bf{r}}x}+\mathfrak{q}_{z}({\bf{r}})\tau_{{\bf{r}}z}\right]$ (3)
where $g$ is the strength of the spin-phonon coupling.
Figure 1: (a) Phase diagram of the pseudospin-1/2 model Hamiltonian from Eqs. 1, 2 and 3 as a function of temperature and $J_{1}/J_{0}$ (see main text for parameters). Green region indicates the strength of the ferro-octupolar order $\langle\tau_{y}\rangle$, and purple region indicating the strength of antiferro-quadrupolar (AFQ) order. White stars are parameter points used in panel (b) and Fig. 2. (b) Phonon dynamics in pump-probe simulations induced by initial momentum kick for $Q_{x}$ phonon, shown for $J_{1}/J_{0}=1.3$ with $T/J_{0}=0.1$ and $\eta/\omega=0.001$ (left: $T<T_{\rm oct}$, single domain) and $T/J_{0}=3.7$ (right: $T>T_{\rm{oct}}$), averaged over 160 different initial conditions. For $T<T_{\rm{oct}}$, we observe coherent energy transfer between $Q_{x}$ and $Q_{z}$ phonons, which is absent for $T>T_{\rm{oct}}$. (c) Comparison showing qualitative agreement between octupolar order parameter $M_{y}=\langle\tau_{y}\rangle$ (blue dots) and $Q_{z}$-$Q_{x}$ cross-amplitude, i.e. maximum $Q_{z}$ response to $Q_{x}$ pump (orange dots) as shown in panel (b); both curves are scaled by their lowest $T$ data point.
Equilibrium phase diagram.— We begin by constructing the equilibrium phase diagram of the full model $H=H_{\rm sp}+H_{\rm ph}+H_{\rm sp-ph}$ in the presence of coupling to the cubic ${\cal E}_{g}$ phonon modes. For this purpose, it is useful to define a length scale $a\!=\!0.01a_{0}$, where $a_{0}\!\approx\!4$ is the cubic lattice constant [70]; in terms of this, a normalized phonon amplitude $Q_{\alpha}=\mathfrak{q}_{\alpha}/a$ of unity corresponds to an experimentally accessible $1\%$ dynamical strain. The corresponding normalized conjugate momenta are $P_{\alpha}=a\mathfrak{p}_{\alpha}$. Measuring energy in units of $J_{0}$ yields dimensionless Hamiltonians,
$H_{\rm sp}$ $=$ $\mathfrak{h}_{\rm sp}/J_{0}$ (4)
$H_{\rm ph}$ $=$ $\!\!\!\sum_{{\bf{r}},\alpha=x,z}\left[\frac{P_{\alpha}^{2}({\bf{r}})}{2M}+\frac{1}{2}M\Omega^{2}Q_{\alpha}^{2}({\bf{r}})\right],$ (5)
$H_{\rm sp-ph}$ $=$ $-\lambda\sum_{\bf{r}}\left[Q_{x}({\bf{r}})\tau_{{\bf{r}}x}+Q_{z}({\bf{r}})\tau_{{\bf{r}}z}\right]$ (6)
where the dimensionless coefficients in the phonon Hamiltonians are
$\frac{1}{2M}=\frac{\hbar^{2}}{2mJ_{0}a^{2}};~{}~{}\Omega=\frac{\hbar\omega}{J_{0}};~{}~{}\lambda=\frac{ga}{J_{0}}$ (7)
This choice of energy units corresponds to measuring time in units of $\hbar/J_{0}$ when we turn to the dynamics. With $m=m_{\rm oxygen}\approx 2.6\times 10^{-26}\rm kg$, and $\hbar\omega\approx 80$ meV as a typical ${\cal E}_{g}$ phonon mode frequency, we find $\frac{1}{2M}\approx 80$, and $\Omega\approx 80$. Density functional theory shows that an approximate $\sim\!1\%$ static strain leads to a splitting $\sim\!5$ meV splitting of the non-Kramers doublet [46], which translates to a dimensionless spin-phonon coupling $\lambda\approx 5$.
Fig. 1(a) displays the phase diagram obtained from our MC simulations (on system size with $10^{3}$ sites) as we vary the temperature $T/J_{0}$ and quadrupolar exchange $J_{1}/J_{0}$. We find a high temperature paramagnetic phase, a low temperature ferro-octupolar ordered phase for $J_{1}/J_{0}\!\lesssim\!1.8$ and an antiferroquadrupolar phase with wavevector $(\pi,\pi,0)$ (and equivalent) for $J_{1}/J_{0}\!\gtrsim\!1.8$. At low $T$, phonon displacements are absent in the ferro-octupolar phase, while the antiferroquadrupolar phase shows small displacements $Q_{\alpha}\!\sim\!0.1$ which track the local quadrupole order. This scale of lattice displacements, $\sim\!0.1\%$, is comparable with those found in the $j=3/2$ Mott insulator Ba2MgReO6 [49, 51, 47]. We next explore the out-of-equilibrium pseudospin dynamics as influenced by pumping and driving of local ${\cal E}_{g}$ phonon modes, with an eye to probe and control octupolar order.
Detection of octupolar order. — In pump-probe experiments [53], an incoming pump beam can excite the even-parity cubic ${\cal E}_{g}$ phonon modes via a two-photon Raman transition or via non-linear phononics, and their subsequent dynamics can potentially provide useful information about broken symmetries in the system. To simulate such pump-probe dynamics, we use equilibriated Monte Carlo configurations of the spins and phonon coordinates as initial conditions, and excite the $Q_{x}$ phonon by choosing its momenta from a displaced Maxwell thermal distribution. In doing so, we model the pump-beam as a spatially uniform impulsive momentum kick to the phonon at all sites [53]; we check that our quench leads to a maximum normalized phonon displacement $Q_{\alpha}\lesssim 1$ over the time evolution. With these initial conditions, we compute the Schrödinger evolution for the pseudospin wavefunction [71] together with Newton’s laws for the classical phonons, and track time evolution of observables averaged over a large number of initial Monte Carlo configurations out to times $t=100\hbar/J_{0}$ for system sizes up to $10^{3}$ sites. We have included a weak damping of the ${\cal E}_{g}$ phonons, incorporating an additional term $(-\eta P_{\alpha})$ in the force equation $dP_{\alpha}/dt$, to capture energy dissipation to the bath of acoustic phonon modes. This mechanism of energy damping via phonons is more relevant to Mott insulators than Gilbert damping of spin modes which plays a role in metallic magnets. Fig. 1(b) shows the time evolution of the phonon displacements $Q_{x}(t)$ and $Q_{z}(t)$ when we give an impulsive kick to the $Q_{x}$ phonon at time $t=0$ in the octupolar ordered and paramagnetic phases. The rapid oscillations of the $Q_{x},Q_{z}$ modes at the phonon frequency are not visible for the long time scales shown in the plot, but are shown in the Supplementary Material (SM) [72]. Remarkably, octupolar broken symmetry for $T<T_{\rm oct}$ mediates coherent energy oscillations between the two phonon modes on time scales $t\sim\!5\hbar/J_{0}$, while such energy oscillations are rapidly dephased or very weak in the paramagnet and quadrupolar phases.
As shown in Fig. 1(c), the maximum cross-amplitude, i.e. the highest amplitude of the $Q_{z}$ oscillations with an impulsive kick to the $Q_{x}$ phonon, serves as a direct probe of the octupolar order parameter. We can understand this with a mean field model of the dynamics (see SM [72]), where the octupolar ordered state leads to an effective field along $\tau_{y}$, and the pumped phonon leads to tipping of the initial pseudospin state away from $\tau_{y}$. This leads to coherent pseudospin precession around the $\tau_{y}$-axis, serving as a conduit to transfer energy from $Q_{x}$ to $Q_{z}$ and back. Thus, by studying the response in the non-pumped phonon mode, it would be possible to identify the presence of octupolar order. Since the phonon modes we consider here are degenerate, one would need a technique to experimentally distinguish between them. This could be achieved using time-resolved X-ray diffraction (trXRD), in which a THz pump pulse initially excites one phonon mode, and X-ray diffraction can be used to measure time-resolved atomic positions and thus extract the amplitudes of the individual modes as a function of time delay [73]. The cross-amplitude of the non-excited phonon mode would be revealed on timescales of the order $t\sim 12\hbar/J_{0}$ (about $8\rm ps$), which is reasonable for trXRD experiments. Using this method, one could distinguish between the different excited phonon modes and thus identify the presence of octupolar order.
Dynamical training of octupolar order. — We next explore how a simultaneous coherent drive of both cubic ${\cal E}_{g}$ phonon modes can be used to control octupolar ordering. We consider a time-dependent Hamiltonian $H_{\rm drive}^{\rm ph}=-A\sum_{\bf{r}}[Q_{x}({\bf{r}})\cos(\Omega t)+Q_{z}({\bf{r}})\cos(\Omega t-\phi)]$ which describes the coherent driving, with drive strength $A$, of the degenerate $Q_{x},Q_{z}$ phonons on resonance, and explore its impact on the pseudospins. Fig. 2(a) shows the effect of this resonant drive in the paramagnetic phase, averaged over 320 initial MC configurations, for phases $\phi=(0,\pm\pi/4,\pm\pi/2)$, where we include weak phonon dissipation to allow the system to thermalize and reach steady state. For a drive strength which leads to a phonon amplitude $Q_{0}\sim 0.5$, we find that, depending on the relative phase, the coherent drive can induce octupolar order $\langle\tau_{y}\rangle$ of either sign in the paramagnetic phase. Previous work has shown, using effective rate equations or local models, that chiral driving of phonons can pump angular momentum and induce effective fields for dipolar $4f$ magnets [23, 24, 25]. Our work provides a nontrivial generalization of this idea to multipolar orders on the lattice. To understand this physics, we note that the phonons linearly couple to the quadrupoles, motivating a proxy Hamiltonian which directly drives quadrupoles as $H^{\rm sp}_{\rm drive}=-A_{\rm eff}\sum_{\bf{r}}(\tau_{{\bf{r}}x}\cos(\Omega t)+\tau_{{\bf{r}}z}\cos(\Omega t-\phi))$. Here, the effective field $A_{\rm eff}=\lambda Q_{0}$ with $Q_{0}$ being the driven phonon amplitude (which depends on the drive strength $A$). We study this Floquet problem using a Magnus expansion (see SM [72] for details) since $\Omega=\hbar\omega/J_{0}\gg 1$, i.e. the phonon dynamics is much faster than the pseudospin dynamics [5, 3, 74]. In standard notation,
$\!H^{\rm sp}_{\rm drive,0}$ $=$ $0;~{}~{}~{}H^{\rm sp}_{\rm drive,\pm}\!=-\frac{\lambda Q_{0}}{2}\!\sum_{\bf{r}}(\tau_{{\bf{r}}x}\!+\!e^{\mp i\phi}\tau_{{\bf{r}}z})$ (8)
$\!\!\!H^{\rm sp,eff}_{\rm drive}\!$ $=$ $\frac{1}{\Omega}[H^{\rm sp}_{\rm drive,+},H^{\rm sp}_{\rm drive,-}]\!=\frac{\lambda^{2}Q_{0}^{2}}{\Omega}\sin\phi\!\sum_{\bf{r}}\tau_{{\bf{r}}y}$ (9)
This phonon-induced effective field in $H^{\rm sp,eff}_{\rm drive}$, $(-\lambda^{2}Q_{0}^{2}/\Omega)\sin\phi$, explains the induced $\phi$-dependent octupolar order for $T\!>\!T_{\rm oct}$ shown in Fig. 2(a).
Figure 2: Dynamics of $\langle\tau_{y}\rangle$ in the presence of phase coherent two-phonon drive with weak dissipation (see text for details). (a) In the paramagnet $T\!=\!1.1T_{\rm oct}$, with drive strength $A\!=\!0.05$ (resonant phonon amplitude $Q_{0}\!\approx\!0.5$) and various phase differences $\phi\in[-\pi/2,\pi/2]$, showing that two-phonon drive can be used ‘train’ the sign of octupolar order. (b) In the ferro-octupolar ordered phase $T\!=\!0.8T_{\rm oct}$, with a larger drive strength $A\!=\!0.2$ (resonant phonon amplitude $Q_{0}\!\approx\!1.5$) showing that with a suitable choice of phase $\phi\!=\!\pi/2$ it can be used to switch from positive to negative octupolar order.
Dynamical switching of octupolar order.— We next ask if, using a strong enough phonon drive, this same mechanism can also be used to switch the sign of octupolar order for $T\!<\!T_{\rm oct}$. Fig. 2(b) shows the impact of the two-phonon coherent drive for $\phi=\pi/2$ at $T/T_{\rm oct}=0.8$. Using a drive which leads to a resonant phonon amplitude $Q_{0}\sim 1.5$, we find that octupolar order in the positive domain switches sign, while the octupolar order in the negative domain is nearly unaffected. The slow evolution of $\langle\tau_{y}\rangle$ at late times, and the nearly unchanged $\langle\tau_{y}\rangle$ in the ‘--’ domain, indicate that heating effects are weak at these timescales. In order for this switching to occur, the phonon-induced field $(-\lambda^{2}Q_{0}^{2}/\Omega)\sin\phi$ must overcome the octupolar Weiss field in the ordered phase. We find that the phonon-induced field is too small to flip pseudospins deep inside a domain, but it is able to induce switching at domain walls where the Weiss field is strongly suppressed. Thus, negative domains must grow by flipping pseudospins at the boundary until the entire octupolar $\langle\tau_{y}\rangle$ flips. In Fig.3, we depict the time evolution of pseudospins under a two-phonon drive starting with an initial 3D domain wall configuration and a thermal population of phonon momenta. We focus on a single 2D slice through the lattice, with colours indicating positive $\langle\tau_{y}\rangle$ (red) and negative $\langle\tau_{y}\rangle$ (blue) and the black sites denoting the non-magnetic sublattice. One can clearly see the two-phonon drive inducing a shrinking of the $+\langle\tau_{y}\rangle$ domain as a function of time via spin flipping at the domain walls. In the bottom right panel of Fig.3 we show the expectation values of $\tau_{x},\tau_{y}$ and $\tau_{z}$ averaged over the lattice; we attribute the rapid oscillations to the high-frequency phonons and the apparent $x$-translational symmetry breaking to the initial distribution of phonon momenta. We provide similar pictures of the lattice spins for the switching presented in Fig. 2(b) in the SM [72], as well as a simple theoretical model of this mechanism.
Figure 3: Visualization of octupolar order switching starting with an initial domain wall configuration of pseudospins. We drive the $Q_{x}$ and $Q_{z}$ phonons with a relative phase of $\pi/2$ and show three time points around $t\sim\!7\hbar/J_{0}$, the point at which switching begins to occur. Here, we depict spins having $\langle\tau_{y}\rangle>0$ as red squares and $\langle\tau_{y}\rangle<0$ as blue squares. We initialize half the spins with $\langle\tau_{y}\rangle=+1.0$ and half with $\langle\tau_{y}\rangle=-1.0$. As time progresses, one can clearly see the shrinking of the positive (red) domain and growth of the negative (blue) domain as reflected in $\langle\tau_{y}\rangle$ averaged over the whole lattice, which is shown in the bottom right panel. Note that the rapid oscillation in the $\tau_{x},\tau_{y}$ and $\tau_{z}$ magnetization appears at the phonon frequency, while the apparent translational symmetry breaking is due to the initial phonon momenta, which are chosen from a thermal distribution.
While experimental verification of our theoretical results on training and switching of octupolar order switching would necessitate a method of detecting the presence of octupolar order, this has proven notoriously difficult to do using conventional experimental techniques. Remarkably, Ref.[75] presents a promising route by which to unambiguously diagnose both the presence and, significantly, sign of the octupolar order parameter. Explicitly, rotational anisotropy second harmonic generation (SHG) measurements are conducted and the presence of octupolar order is characterized through interference between the time reversal even and odd channels, with the sign of the octupolar order determining the positions of the measured intensity maxima. Thus, we propose a tr-SHG experiment with a first measurement conducted at $t=0$ and a second conducted at time $t\sim 25\hbar/J_{0}$ ($\sim 16\rm ps$) looking for changes in the position of intensity maxima as a potential route to diagnose the training and switching of the octupolar order parameter. In the SM [72], we also provide further plots showing how our results depend on the phenomenological damping parameter $\eta$.
Summary.— We have employed Monte Carlo and molecular dynamics simulations capable of modelling lattice degrees of freedom, in the form of local Einstein phonons, in addition to multipolar degrees of freedom to study Mott insulators which host non-Kramers doublets. Our equilibrium phase diagram for the model hosts ferro-octupolar or antiferro-quadrupolar ordered phases at low temperature, as partly described by previous work, but shows that the quadrupolar order is accompanied by lattice displacements. We next studied the out-of-equilibrium multipolar-phonon dynamics in the presence of a uniform excitation of Einstein phonon modes. Emulating pump-probe type experiments by giving a uniform impulsive kick to one phonon mode, we observed a characteristic response in the non-excited phonon below the octupolar ordering temperature which is absent in the disordered phase, leading to a dynamical signature of equilibrium octupolar order. Secondly, we model a continuous excitation of both phonons, with a relative phase between them, and observe the generation of octupolar order consistent with an effective field predicted by a lab frame Magnus expansion. This also led us to a proposed route to ultrafast picosecond timescale switching of octupolar order using phonons. Our work opens up unique directions in the study of Mott insulators playing host to multipolar moments, including potential routes by which to detect and control complex hidden orders.
This research was funded by an NSERC Discovery Grant (AP), an Ontario Graduate Scholarship (KH), and an NSERC CGS-D fellowship (RS). We thank Sreekar Voleti for initial collaboration on this project. We acknowledge Dave Hsieh, Emily Zhang, William Bateman-Hemphill, and Cristian Batista for extremely helpful discussions. Numerical computations were performed on the Niagara supercomputer at the SciNet HPC Consortium and the Digital Research Alliance of Canada.

### v4

@@
 As a case study, we examine pseudospin-1/2 Mott insulators hosting quadrupolar and octupolar moments, and investigate the effect of resonantly exciting ${\cal E}_{g}$ phonon modes which couple linearly to the quadrupoles.
-We show that this leads to multipolar precession, with the back action resulting in chiral phonon dynamics below the octupolar phase transition.
-We further explore the impact of a coherent two-phonon Floquet drive, showing that it can be used to ‘train’ octupolar order or even switch its sign on ultrafast timescales.
-Our results are obtained combining a generalized Monte Carlo code incorporating phonons with molecular dynamics simulations which numerically integrate the coupled spin-phonon equations of motion.
-Our work shows how phonon-driven dynamics can be used to probe and control hidden orders in solids.
+We show that this leads to multipolar precession, with the backaction resulting in pseudochiral phonon dynamics in the octupolar ordered phase.
+We further study the impact of a short-time two-phonon drive, showing that it can induce the build-up of octupolar order or even switch its sign on picosecond timescales.
+Our results are obtained using a Monte Carlo code incorporating phonons, molecular dynamics simulations to numerically integrate the coupled spin-phonon equations of motion, and analytical Floquet theory.
+Our work shows how driven phonons can probe and control hidden orders in solids.
 preprint: APS/123-QED
-Introduction.— Shaking electrons or atoms with light has been proposed to induce remarkable non-equilibrium many body states in solid state crystals as well as ultracold atoms [1, 2, 3, 4, 5, 6, 7].
-In recent years, rapid advances in ultrafast and THz optics have led to the ability to resonantly excite specific orbital transitions and phonon modes in solids [8, 9, 10], providing a distinct pathway to control broken symmetry states including ferroelectricity [11], superconductivity [12], and magnetism [13, 9, 14, 15].
-In dipolar magnets, local Einstein phonons can strongly impact the single-ion anisotropy or spin-bilinear exchange interaction on bonds.
-This has been shown to stabilize unconventional equilibrium magnetic orders in geometrically frustrated lattices such as kagome, triangular, or pyrochlore systems [16, 17, 18, 19].
-Going beyond equilibrium orders, recent experimental and theoretical work has shown that coherent phonon driving can produce fascinating non-equilibrium Floquet dynamics in spin-phonon coupled dipolar magnets.
-These include the generation of anomalously large effective magnetic fields from driven chiral phonons [20, 21, 22, 23, 24, 25, 26], tuning of magnetic orders and magnon band topology [27, 9, 14, 28], frustration breaking in dimerized magnets [29], nonequilibrium phase transitions [30], and high harmonic generation [31].
-These advances have motivated the study of many-body coupled spin-phonon dynamics in such systems while incorporating dissipation effects.
-Earlier work has shown the value of such an approach for understanding pump-probe dynamics in layered van der Waals magnets [32].
-Going beyond the physics of dipolar spin systems, there is a rich set of issues one encounters when one starts to explore multipolar degrees of freedom which are found in solids with strong spin-orbit coupling.
+Introduction.— Shaking electrons or atoms with light can induce remarkable non-equilibrium many body states in solid state crystals as well as ultracold atoms [1, 2, 3, 4, 5, 6, 7, 8].
+In recent years, rapid advances in ultrafast and THz optics have led to the ability to resonantly excite specific orbital transitions and phonon modes in solids [9, 10, 11, 12], providing a distinct pathway to control broken symmetry states including ferroelectricity [13], superconductivity [14], and magnetism [15, 10, 16, 17, 18, 19].
+In dipolar magnets, Einstein phonons can strongly impact the spin-spin exchange interaction, stabilizing unconventional equilibrium magnetic orders in geometrically frustrated lattices such as kagome, triangular, or pyrochlore systems [20, 21, 22, 23].
+In addition to equilibrium orders, recent experimental and theoretical work has shown that coherent phonon driving can produce fascinating non-equilibrium Floquet dynamics in spin-phonon coupled dipolar magnets.
+These include anomalously large effective magnetic fields from driven chiral phonons [24, 25, 26, 27, 28, 29, 30], tuning of magnetic orders and magnon band topology [31, 10, 16, 32], frustration breaking in dimerized magnets [33], nonequilibrium phase transitions [34], and high harmonic generation [35].
+In addition, this has motivated the study of many-body spin-phonon dynamics and Floquet-phonon physics in layered van der Waals materials [36, 37, 38, 39].
+Going beyond dipolar spin systems, there is a rich set of issues one encounters for multipolar degrees of freedom which are found in solids with strong spin-orbit coupling.
 Unlike dipole moments, higher order multipoles such as quadrupoles and octupoles are associated with spin-orbital entangled charge or current distributions which do not simply couple to conventional probes; for this reason, multipolar broken symmetries, found in diverse materials, are hard to detect and generically termed ‘hidden orders’ [40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53].
-Quadrupolar moments, however, being even under time-reversal symmetry, can linearly couple to phonons, which can lead to a significant differences in the equilibrium ordering and excitations of multipoles when compared with magnetic dipoles.
-This is exemplified by the recent understanding of unusual patterns of multipolar symmetry breaking and lattice displacements in Ba2MgReO6 [47, 48, 49, 50, 51] and the direct detection of vibronic modes by observing spectral splitting and side-bands in resonant inelastic X-ray scattering in K2IrCl6 [52].
+Quadrupolar moments, however, being even under time-reversal symmetry, can linearly couple to phonons, leading to significant differences in the equilibrium ordering and excitations of multipoles when compared with magnetic dipoles.
+This is exemplified by recent discoveries of unusual multipolar symmetry breaking and lattice displacements in Ba2MgReO6 [54, 55, 56, 57, 58] and the detection of vibronic modes via resonant inelastic X-ray scattering in K2IrCl6 [59].
 Recent experiments have begun to extend optical and THz techniques to study the non-equilibrium dynamics of multipolar orders.
-For instance, resonantly excited phonon modes, on ultrafast timescales, in pump-probe and THz experiments have shown evidence of multipolar modes coupled to lattice vibrations [29].
+For instance, resonantly excited phonon modes in pump-probe and THz experiments have shown evidence of multipolar modes coupled to lattice vibrations [33].
 In $\rm{Ca_{2}RuO_{4}}$, THz experiments have revealed a ‘hidden’ quadrupolar-ordered state through the anomalous broadening of coherently driven phonons [60].
-Such coupling between phonons and quadrupole moments suggests that it may be possible to not only probe but also achieve the dynamical control of multipolar orders using optical and THz techniques.
-These developments provide an impetus for the theoretical exploration of pumped and driven multipolar materials, and call out for a theory of many-body dynamics in coupled multipolar-phonon systems.
-In this Letter, we fill this important gap by developing a simulation tool to study multipolar orders and their semiclassical dynamics including the crucial effect of coupling to local Einstein phonons and incorporating dissipation effects.
-In order to do this, we combine large scale equilibrium spin-phonon Monte Carlo (MC) simulations with spin-phonon molecular dynamics (MD) obtained by integrating their coupled equations of motion.
-The key conceptual idea underlying our work is that a phonon mode coupled to a quadrupole operator acts as a transverse field to all other multipole operators which do not commute with the relevant quadrupole.
-Kicking the phonon thus engenders nontrivial precessional dynamics in multipolar space which can be studied theoretically and experimentally, enabling one to shed light on complex hidden orders.
-While our approach is broadly applicable to many multipolar materials, we illustrate the power of this tool by focussing on a recently discovered paradigmatic example of multipolar systems in Mott insulators, the osmate double perovskites [54, 55, 56, 46, 57, 58] which have been recently shown to display competing octupolar and quadrupolar ordering tendencies [55, 56, 58].
-Our work shows how the presence of local phonons, resonantly excited by either pump pulses or Floquet driving, can influence the dynamics of the multipolar moments in such a way as to enable the detection, training, and ultrafast switching of exotic octupolar order.
-Our results point to a number of future research directions, including experiments to probe and control elusive hidden orders using phonons.
+These developments call out for a theory of many-body dynamics in pumped and driven multipolar quantum materials as a route to probe and control hidden orders.
+In this Letter, we fill this important gap by developing a simulation tool to study multipolar orders and their semiclassical dynamics including the crucial effect of coupling to Einstein phonons with dissipation.
+This method combines large scale equilibrium spin-phonon Monte Carlo (MC) simulations with spin-phonon molecular dynamics (MD) obtained by integrating their coupled equations of motion.
+The key idea underlying our work is that a phonon mode coupled to a quadrupole acts as a transverse field to all other orthogonal multipoles, just as for a dipolar spin, a magnetic field along $S_{z}$ acts as a transverse field to $(S_{x},S_{y})$.
+For multipolar materials, kicking the phonon thus engenders nontrivial precessional dynamics in multipolar space which we leverage as a route to unearth hidden orders.
+While our approach is broadly applicable to many multipolar materials, we illustrate the power of this tool by focusing on a recently discovered paradigmatic example in Mott insulators, the osmate double perovskites [61, 62, 63, 53, 64, 65] which display competing octupolar and quadrupolar ordering tendencies [62, 63, 65].
+Our work shows how the presence of phonons, resonantly excited by either pump pulses or Floquet driving, can influence the dynamics of the multipolar moments and enable the detection, build-up, and ultrafast switching of exotic octupolar order.
 Model.— The simplest example of multipolar moments arise in materials with non-Kramers doublets, local pseudospin-1/2 degrees of freedom whose doublet degeneracy is protected by crystalline point group symmetries rather than just time-reversal symmetry as in the case of Kramers doublets.
-The ability to split this doublet by breaking the local point group symmetry shows that these doublets have quadrupolar character and are not pure dipoles.
 Remarkably, non-Kramers doublets are found in a diverse array of quantum materials including: quantum spin-ice compounds such as Pr2Zr2O7, Pr2Sn2O7 [66, 67, 68, 69], Kondo lattice intermetallics PrTi2Al20, PrV2Al20 [70, 71, 49, 72, 50], Mott insulating double perovskites $\rm{Ba_{2}CaOsO_{6}}$, $\rm{Ba_{2}MgOsO_{6}}$, $\rm{Ba_{2}ZnOsO_{6}}$ [61, 62, 53], and vacancy ordered double perovskites Cs2WCl6 and Rb2WCl6 [73, 74].
 We focus here on non-Kramers doublets with time-reversal even (${\cal T}$-even) quadrupolar moments represented by pseudospin-1/2 Pauli operators $(\tau_{x},\tau_{z})$ and a ${\cal T}$-odd octupole represented by $\tau_{y}$ as realized in the double perovskites and the Kondo intermetallics.
-In the double perovskite Mott insulators, these pseudospins live on a face-centered cubic (fcc) lattice, and descend from the splitting of a $J=2$ angular momentum multiplet, with $\tau_{x}\propto(J_{x}^{2}-J_{y}^{2})$, $\tau_{z}\propto(3J_{z}^{2}-J(J+1))$, and $\tau_{y}\propto\overline{J_{x}J_{y}J_{z}}$, with bar denoting operator symmetrization.
+In the double perovskites, these pseudospins live on a face-centered cubic (fcc) lattice (see End Matter), and descend from splitting a $J=2$ angular momentum multiplet, with $\tau_{x}\propto(J_{x}^{2}-J_{y}^{2})$, $\tau_{z}\propto(3J_{z}^{2}-J(J+1))$, and $\tau_{y}\propto\overline{J_{x}J_{y}J_{z}}$, with bar denoting operator symmetrization.
 Symmetries dictate the following fcc lattice Hamiltonian describing nearest-neighbor pseudospin interactions
@@
 The phase diagram of this model depends on the dimensionless ratios of exchange couplings via $J_{1}/J_{0}$ and $\gamma$ [75, 63].
-Based on density functional theory and dynamical mean field theory for Ba2CoOsO6 [69], we set $\gamma\!\approx\!-0.4$ and $J_{1}/J_{0}\!\approx\!0.5$ for the pseudospin Hamiltonian; this places us in the ferro-octupolar phase for $T/J_{0}\lesssim 3.2$ as determined from classical Monte Carlo simulations [56].
-A microscopic calculation of the octupolar exchange yields $J_{0}\approx 1$ meV, which leads to $T_{\rm oct}\approx 40$ K, in good agreement with experimental results on Ba2CoOsO6 [54, 56, 46].
-Motivated by constructing the more general phase diagram of this model, we fix $\gamma$ and vary $J_{1}/J_{0}$ and $T/J_{0}$ which allows us to access both ferro-octupolar and quadrupolar ordered phases.
+Based on density functional theory (DFT) and dynamical mean field theory for Ba2CoOsO6 [76], we set $\gamma\!\approx\!-0.4$ and $J_{1}/J_{0}\!\approx\!0.5$ for the pseudospin Hamiltonian; this places us in the ferro-octupolar phase for $T/J_{0}\lesssim 3.2$ as determined from classical Monte Carlo simulations [63].
+A microscopic calculation of the octupolar exchange yields $J_{0}\approx 1$ meV, which leads to $T_{\rm oct}\approx 40$ K, in good agreement with experiments on Ba2CoOsO6 [61, 63, 53].
+To construct a general phase diagram, we fix $\gamma$ and vary $J_{1}/J_{0}$ and $T/J_{0}$ to access both ferro-octupolar and quadrupolar ordered phases.
 Next, we consider cubic symmetry ${\cal E}_{g}$ Einstein phonon modes at each lattice site with $x\equiv x^{2}\!-\!y^{2}$, $z\equiv 3z^{2}\!-\!r^{2}$, with the phonon Hamiltonian
@@
 where $g$ is the strength of the spin-phonon coupling.
-Figure 1: (a) Phase diagram of the pseudospin-1/2 model Hamiltonian from Eqs.
-1, 2 and 3 as a function of temperature and $J_{1}/J_{0}$ (see main text for parameters).
-Green region indicates the strength of the ferro-octupolar order $\langle\tau_{y}\rangle$, and purple region indicating the strength of antiferro-quadrupolar (AFQ) order.
-White stars are parameter points used in panel (b) and Fig.
-2.
-(b) Phonon dynamics in pump-probe simulations induced by initial momentum kick for $Q_{x}$ phonon, shown for $J_{1}/J_{0}=1.3$ with $T/J_{0}=0.1$ and $\eta/\omega=0.001$ (left: $T<T_{\rm oct}$, single domain) and $T/J_{0}=3.7$ (right: $T>T_{\rm{oct}}$), averaged over 160 different initial conditions.
+Figure 1: (a) Monte Carlo phase diagram of the Hamiltonian in Eqs.
+1, 2 and 3 as a function of $T/J_{0}$ and $J_{1}/J_{0}$ (for fixed $\gamma=-0.4$).
+Colors in the ordered phases depict strength of the ferro-octupolar order $\langle\tau_{y}\rangle$ (green) and antiferro-quadrupolar (AFQ) order (purple); insets depict the AFQ and ferro-octupolar orders on two adjacent layers of the fcc lattice (top layer: bold/large, bottom layer: light/small) with orbital shapes and colors respectively showing the charge density and the sign, and the ‘+’ for AFQ indicating a superposition of the two orbital orders.
+(b) and (c) Phonon dynamics in pump-probe simulations induced by initial momentum kick for $Q_{x}$ phonon for $J_{1}/J_{0}=1.3$ and phonon damping constant $0.072\pi\eta/\omega=$ 0.07 [77, 78], averaged over 160 different initial conditions.
+(b) $T=0.1J_{0}<T_{\rm oct}$, single domain ferro-octupolar; (c) $T=5.2J_{0}>T_{\rm{oct}}$, paramagnet.
 For $T<T_{\rm{oct}}$, we observe coherent energy transfer between $Q_{x}$ and $Q_{z}$ phonons, which is absent for $T>T_{\rm{oct}}$.
-(c) Comparison showing qualitative agreement between octupolar order parameter $M_{y}=\langle\tau_{y}\rangle$ (blue dots) and $Q_{z}$-$Q_{x}$ cross-amplitude, i.e.
-maximum $Q_{z}$ response to $Q_{x}$ pump (orange dots) as shown in panel (b); both curves are scaled by their lowest $T$ data point.
+(d) Comparison showing qualitative agreement between octupolar order parameter $M_{y}=\langle\tau_{y}\rangle$ (blue dots) and $Q_{z}$-$Q_{x}$ cross-amplitude, i.e.
+maximum $Q_{z}$ response to $Q_{x}$ pump (orange dots) as shown in panels (b) and (c); both curves are scaled by their lowest $T$ data point.
+(e) Schematic of energy exchange between the $(Q_{x},Q_{z})$ phonons in the ferro-octupolar phase through pseudospin precession and quadrupole-phonon coupling.
 Equilibrium phase diagram.— We begin by constructing the equilibrium phase diagram of the full model $H=H_{\rm sp}+H_{\rm ph}+H_{\rm sp-ph}$ in the presence of coupling to the cubic ${\cal E}_{g}$ phonon modes.
-For this purpose, it is useful to define a length scale $a\!=\!0.01a_{0}$, where $a_{0}\!\approx\!4$ is the cubic lattice constant [70]; in terms of this, a normalized phonon amplitude $Q_{\alpha}=\mathfrak{q}_{\alpha}/a$ of unity corresponds to an experimentally accessible $1\%$ dynamical strain.
-The corresponding normalized conjugate momenta are $P_{\alpha}=a\mathfrak{p}_{\alpha}$.
+For this purpose, it is useful to define a length scale $a\!=\!0.01a_{0}$, where $a_{0}\!\approx\!4$ is the cubic lattice constant [79], and a normalized phonon amplitude $Q_{\alpha}=\mathfrak{q}_{\alpha}/a$; in these units, $Q_{\alpha}=1$ implies an experimentally accessible $1\%$ atomic displacement, relative to the lattice constant [15].
+The normalized conjugate momenta are $P_{\alpha}=a\mathfrak{p}_{\alpha}$.
 Measuring energy in units of $J_{0}$ yields dimensionless Hamiltonians,
@@
 $H_{\rm sp-ph}$ $=$ $-\lambda\sum_{\bf{r}}\left[Q_{x}({\bf{r}})\tau_{{\bf{r}}x}+Q_{z}({\bf{r}})\tau_{{\bf{r}}z}\right]$ (6)
-where the dimensionless coefficients in the phonon Hamiltonians are
-$\frac{1}{2M}=\frac{\hbar^{2}}{2mJ_{0}a^{2}};~{}~{}\Omega=\frac{\hbar\omega}{J_{0}};~{}~{}\lambda=\frac{ga}{J_{0}}$ (7)
-This choice of energy units corresponds to measuring time in units of $\hbar/J_{0}$ when we turn to the dynamics.
-With $m=m_{\rm oxygen}\approx 2.6\times 10^{-26}\rm kg$, and $\hbar\omega\approx 80$ meV as a typical ${\cal E}_{g}$ phonon mode frequency, we find $\frac{1}{2M}\approx 80$, and $\Omega\approx 80$.
-Density functional theory shows that an approximate $\sim\!1\%$ static strain leads to a splitting $\sim\!5$ meV splitting of the non-Kramers doublet [46], which translates to a dimensionless spin-phonon coupling $\lambda\approx 5$.
+where the dimensionless coefficients in Eqn.
+5, 6 are $1/2M=\hbar^{2}/2mJ_{0}a^{2}$, $\Omega=\hbar\omega/J_{0}$, and $\lambda=ga/J_{0}$.
+This choice of energy units corresponds to measuring time in units of $\hbar/J_{0}$ ($\approx\!0.6$ picosecond).
+We set $m=m_{\rm oxygen}\approx 2.6\times 10^{-26}\rm kg$, since we are simulating ${\cal E}_{g}$ phonon modes where only the oxygen atoms have finite displacements, and we set $\hbar\omega\approx 80$ meV as a typical ${\cal E}_{g}$ phonon mode frequency; using this, $1/2M\approx 80$, and $\Omega\approx 80$.
+From DFT, a $\sim\!1\%$ atomic displacement leads to a $\sim\!5$ meV splitting of the non-Kramers doublet [53]; this implies $g=5{\rm meV}/a$, so the dimensionless spin-phonon coupling $\lambda\approx 5$.
 Fig.
-1(a) displays the phase diagram obtained from our MC simulations (on system size with $10^{3}$ sites) as we vary the temperature $T/J_{0}$ and quadrupolar exchange $J_{1}/J_{0}$.
-We find a high temperature paramagnetic phase, a low temperature ferro-octupolar ordered phase for $J_{1}/J_{0}\!\lesssim\!1.8$ and an antiferroquadrupolar phase with wavevector $(\pi,\pi,0)$ (and equivalent) for $J_{1}/J_{0}\!\gtrsim\!1.8$.
-At low $T$, phonon displacements are absent in the ferro-octupolar phase, while the antiferroquadrupolar phase shows small displacements $Q_{\alpha}\!\sim\!0.1$ which track the local quadrupole order.
-This scale of lattice displacements, $\sim\!0.1\%$, is comparable with those found in the $j=3/2$ Mott insulator Ba2MgReO6 [49, 51, 47].
-We next explore the out-of-equilibrium pseudospin dynamics as influenced by pumping and driving of local ${\cal E}_{g}$ phonon modes, with an eye to probe and control octupolar order.
+1(a) displays the phase diagram obtained from our MC simulations (on system size with $10^{3}$ sites) as we vary the temperature $T/J_{0}$ and quadrupolar exchange $J_{1}/J_{0}$, effectively sampling spin and phonon configurations from thermal populations.
+We find a high temperature paramagnetic phase, a low temperature ferro-octupolar phase for $J_{1}/J_{0}\!\lesssim\!1.8$ and an antiferro-quadrupolar (AFQ) phase with wavevector $(\pi,\pi,0)$ (and equivalent) for $J_{1}/J_{0}\!\gtrsim\!1.8$.
+At low $T$, phonon displacements are absent in the ferro-octupolar phase, while the AFQ phase shows small displacements $Q_{\alpha}\!\sim\!0.1$ which track the local quadrupole order.
+While $Q_{\alpha}({\bf{r}})\neq 0$ at each site, the sign of $Q_{\alpha}({\bf{r}})$ alternates from layer to layer (see Fig.
+1).
+Comparable staggered lattice displacements, $\sim\!0.1\%$, have been found in the AFQ phase of the $j=3/2$ Mott insulator Ba2MgReO6 [56, 58, 54].
+We next explore the out-of-equilibrium pseudospin-phonon dynamics.
 Detection of octupolar order.
-— In pump-probe experiments [53], an incoming pump beam can excite the even-parity cubic ${\cal E}_{g}$ phonon modes via a two-photon Raman transition or via non-linear phononics, and their subsequent dynamics can potentially provide useful information about broken symmetries in the system.
-To simulate such pump-probe dynamics, we use equilibriated Monte Carlo configurations of the spins and phonon coordinates as initial conditions, and excite the $Q_{x}$ phonon by choosing its momenta from a displaced Maxwell thermal distribution.
-In doing so, we model the pump-beam as a spatially uniform impulsive momentum kick to the phonon at all sites [53]; we check that our quench leads to a maximum normalized phonon displacement $Q_{\alpha}\lesssim 1$ over the time evolution.
-With these initial conditions, we compute the Schrödinger evolution for the pseudospin wavefunction [71] together with Newton’s laws for the classical phonons, and track time evolution of observables averaged over a large number of initial Monte Carlo configurations out to times $t=100\hbar/J_{0}$ for system sizes up to $10^{3}$ sites.
-We have included a weak damping of the ${\cal E}_{g}$ phonons, incorporating an additional term $(-\eta P_{\alpha})$ in the force equation $dP_{\alpha}/dt$, to capture energy dissipation to the bath of acoustic phonon modes.
-This mechanism of energy damping via phonons is more relevant to Mott insulators than Gilbert damping of spin modes which plays a role in metallic magnets.
+— In pump-probe experiments [60], an incoming pump beam can excite the even-parity cubic ${\cal E}_{g}$ phonon modes via a two-photon Raman transition or via non-linear phononics, with the subsequent dynamics providing useful information about broken symmetries.
+To simulate pump-probe dynamics, we use equilibriated MC spin-phonon configurations as initial conditions, impulsively excite the $Q_{x}$ phonon, and time-evolve the coupled spin and phonon degrees of freedom; see End Matter for simulation details.
 Fig.
-1(b) shows the time evolution of the phonon displacements $Q_{x}(t)$ and $Q_{z}(t)$ when we give an impulsive kick to the $Q_{x}$ phonon at time $t=0$ in the octupolar ordered and paramagnetic phases.
-The rapid oscillations of the $Q_{x},Q_{z}$ modes at the phonon frequency are not visible for the long time scales shown in the plot, but are shown in the Supplementary Material (SM) [72].
-Remarkably, octupolar broken symmetry for $T<T_{\rm oct}$ mediates coherent energy oscillations between the two phonon modes on time scales $t\sim\!5\hbar/J_{0}$, while such energy oscillations are rapidly dephased or very weak in the paramagnet and quadrupolar phases.
+1(b) shows the time evolution of the phonon displacements $Q_{x}(t)$ and $Q_{z}(t)$ when we give an impulsive kick to the $Q_{x}$ phonon at time $t=0$ in the ferro-octupolar and paramagnetic phases.
+The rapid oscillations of the $Q_{x},Q_{z}$ modes at the phonon frequency are not visible in the plot, but are shown in the Supplementary Material (SM) [80].
+Remarkably, we see that octupolar broken symmetry for $T<T_{\rm oct}$ mediates coherent energy transfer from the $Q_{x}$ to $Q_{z}$ mode on a time scale $t\!\approx\!5\hbar/J_{0}$ ($\approx 3$ ps).
+Such energy transfer is absent due to rapid dephasing in the paramagnet ($T\!>\!T_{\rm oct}$) or in the quadrupolar phase; this is shown in Fig.
+1(c) for the paramagnet.
 As shown in Fig.
-1(c), the maximum cross-amplitude, i.e.
-the highest amplitude of the $Q_{z}$ oscillations with an impulsive kick to the $Q_{x}$ phonon, serves as a direct probe of the octupolar order parameter.
-We can understand this with a mean field model of the dynamics (see SM [72]), where the octupolar ordered state leads to an effective field along $\tau_{y}$, and the pumped phonon leads to tipping of the initial pseudospin state away from $\tau_{y}$.
-This leads to coherent pseudospin precession around the $\tau_{y}$-axis, serving as a conduit to transfer energy from $Q_{x}$ to $Q_{z}$ and back.
-Thus, by studying the response in the non-pumped phonon mode, it would be possible to identify the presence of octupolar order.
-Since the phonon modes we consider here are degenerate, one would need a technique to experimentally distinguish between them.
-This could be achieved using time-resolved X-ray diffraction (trXRD), in which a THz pump pulse initially excites one phonon mode, and X-ray diffraction can be used to measure time-resolved atomic positions and thus extract the amplitudes of the individual modes as a function of time delay [73].
-The cross-amplitude of the non-excited phonon mode would be revealed on timescales of the order $t\sim 12\hbar/J_{0}$ (about $8\rm ps$), which is reasonable for trXRD experiments.
-Using this method, one could distinguish between the different excited phonon modes and thus identify the presence of octupolar order.
-Dynamical training of octupolar order.
-— We next explore how a simultaneous coherent drive of both cubic ${\cal E}_{g}$ phonon modes can be used to control octupolar ordering.
-We consider a time-dependent Hamiltonian $H_{\rm drive}^{\rm ph}=-A\sum_{\bf{r}}[Q_{x}({\bf{r}})\cos(\Omega t)+Q_{z}({\bf{r}})\cos(\Omega t-\phi)]$ which describes the coherent driving, with drive strength $A$, of the degenerate $Q_{x},Q_{z}$ phonons on resonance, and explore its impact on the pseudospins.
+1(d), the maximum cross-amplitude, i.e.
+the highest amplitude of $Q_{z}$ oscillations for an impulsive kick to the $Q_{x}$ phonon, serves as a direct probe of the octupolar order parameter.
+We understand this as follows (see also SM [80]).
+The octupolar order leads to an average internal field along $\tau_{y}$, while the pumped phonon tips the initial pseudospin state away from $\tau_{y}$.
+This leads to coherent pseudospin precession around the $\tau_{y}$-axis, serving as a conduit to transfer energy from $Q_{x}$ to $Q_{z}$ and back; see Fig.
+1(e).
+This protocol could be realized using a stimulated Raman scattering mechanism [81] to pump one phonon mode followed by time-resolved X-ray diffraction (trXRD) to probe time-dependent atomic positions for extracting the amplitudes of both phonon modes as a function of time delay.
+Ref.
+[82] identifies deformations of order $\sim 3\times 10^{-2}$Å for a bond of length $\sim 2.67$Å, making feasible the detection of the $\sim 1\%$ distortions discussed in our work.
+The maximum amplitude of the $Q_{z}$ phonon mode, a signature of octupolar order, is revealed on timescales $t\sim 2\hbar/J_{0}$ (about $1.3\rm ps$), which is reasonable for trXRD experiments.
+Dynamical build-up of octupolar order.
+— We next explore how a simultaneous drive of both ${\cal E}_{g}$ phonon modes can be used to control octupolar ordering.
+We consider $H_{\rm drive}^{\rm ph}\!=\!-A(t)\sum_{\bf{r}}[Q_{x}({\bf{r}})\cos(\Omega t)\!+\!Q_{z}({\bf{r}})\cos(\Omega t-\phi)]$ which describes a coherent drive of Raman-active phonons as a result of terahertz sum-frequency excitation; the drive strength $A(t)$ is thus proportional to the square of the electric field [77, 78].
+We implement a short-pulse drive amplitude $A(t)\!=\!A$ for $t\!<\!3\hbar/J_{0}$ and $A(t)\!=\!0$ otherwise.
 Fig.
-2(a) shows the effect of this resonant drive in the paramagnetic phase, averaged over 320 initial MC configurations, for phases $\phi=(0,\pm\pi/4,\pm\pi/2)$, where we include weak phonon dissipation to allow the system to thermalize and reach steady state.
-For a drive strength which leads to a phonon amplitude $Q_{0}\sim 0.5$, we find that, depending on the relative phase, the coherent drive can induce octupolar order $\langle\tau_{y}\rangle$ of either sign in the paramagnetic phase.
-Previous work has shown, using effective rate equations or local models, that chiral driving of phonons can pump angular momentum and induce effective fields for dipolar $4f$ magnets [23, 24, 25].
+2(a) shows the effect of this drive in the paramagnetic phase, averaged over 320 initial MC configurations, for phases $\phi=(0,\pm\pi/4,\pm\pi/2)$.
+For a drive strength which leads to a peak phonon amplitude $Q_{0}\sim 0.75$, we find that, depending on the relative phase, the drive pulse can induce transient octupolar order $\langle\tau_{y}\rangle$ of either sign, which vanishes at long times.
+Previous work has shown, using effective rate equations for single-site models, that chiral driving of phonons can induce effective fields for dipolar $4f$ magnets [27, 83, 28].
 Our work provides a nontrivial generalization of this idea to multipolar orders on the lattice.
-To understand this physics, we note that the phonons linearly couple to the quadrupoles, motivating a proxy Hamiltonian which directly drives quadrupoles as $H^{\rm sp}_{\rm drive}=-A_{\rm eff}\sum_{\bf{r}}(\tau_{{\bf{r}}x}\cos(\Omega t)+\tau_{{\bf{r}}z}\cos(\Omega t-\phi))$.
-Here, the effective field $A_{\rm eff}=\lambda Q_{0}$ with $Q_{0}$ being the driven phonon amplitude (which depends on the drive strength $A$).
-We study this Floquet problem using a Magnus expansion (see SM [72] for details) since $\Omega=\hbar\omega/J_{0}\gg 1$, i.e.
-the phonon dynamics is much faster than the pseudospin dynamics [5, 3, 74].
+To understand this physics, we note that the phonons couple linearly to the quadrupoles, motivating a proxy Hamiltonian $H^{\rm sp}_{\rm drive}=-A_{\rm eff}\sum_{\bf{r}}(\tau_{{\bf{r}}x}\cos(\Omega t)+\tau_{{\bf{r}}z}\cos(\Omega t-\phi))$ which directly drives quadrupoles.
+The effective field $A_{\rm eff}=\lambda Q_{0}$ with $Q_{0}$ being the peak phonon amplitude (which depends on the drive strength $A$).
+We study this using a Floquet-Magnus expansion [80] since $\Omega=\hbar\omega/J_{0}\gg 1$, i.e.
+the phonon dynamics is much faster than pseudospins [5, 3, 84].
 In standard notation,
@@
 2(a).
-Figure 2: Dynamics of $\langle\tau_{y}\rangle$ in the presence of phase coherent two-phonon drive with weak dissipation (see text for details).
-(a) In the paramagnet $T\!=\!1.1T_{\rm oct}$, with drive strength $A\!=\!0.05$ (resonant phonon amplitude $Q_{0}\!\approx\!0.5$) and various phase differences $\phi\in[-\pi/2,\pi/2]$, showing that two-phonon drive can be used ‘train’ the sign of octupolar order.
-(b) In the ferro-octupolar ordered phase $T\!=\!0.8T_{\rm oct}$, with a larger drive strength $A\!=\!0.2$ (resonant phonon amplitude $Q_{0}\!\approx\!1.5$) showing that with a suitable choice of phase $\phi\!=\!\pi/2$ it can be used to switch from positive to negative octupolar order.
-Dynamical switching of octupolar order.— We next ask if, using a strong enough phonon drive, this same mechanism can also be used to switch the sign of octupolar order for $T\!<\!T_{\rm oct}$.
+Figure 2: Dynamics of $\langle\tau_{y}\rangle$ in the presence of short-time phase coherent two-mode phonon drive with weak dissipation.
+The drive $H_{\rm drive}^{\rm ph}$ is turned on from $t=0\hbar/J_{0}$ to $t=3\hbar/J_{0}$ and then turned off.
+(a) In the paramagnet $T\!=\!1.1T_{\rm oct}$, with drive strength $A\!=\!0.5$ (peak phonon amplitude $Q_{0}\!\approx\!0.75$) and various phase differences $\phi\in[-\pi/2,\pi/2]$, showing that driven phonons can select the sign of generated octupolar order.
+(b) In the ferro-octupolar ordered phase $T\!=\!0.8T_{\rm oct}$, with varying drive strength $A$, which induces a peak phonon amplitude $Q_{0}$ between $2-3$.
+With a suitable phase choice $\phi\!=\!\pi/2$, it can be used to decrease the octupolar order parameter $\langle\tau_{y}\rangle$, and even switch from positive to negative.
+Inset: Phonon coordinate $Q_{x}(t)$ for $A\!=\!2$, showing its amplitude build-up and eventual dissipation after the drive is switched off.
+Dynamical switching of octupolar order.— We next ask if this pulsed phonon drive can also be used to switch the sign of octupolar order for $T\!<\!T_{\rm oct}$.
 Fig.
 2(b) shows the impact of the two-phonon coherent drive for $\phi=\pi/2$ at $T/T_{\rm oct}=0.8$.
-Using a drive which leads to a resonant phonon amplitude $Q_{0}\sim 1.5$, we find that octupolar order in the positive domain switches sign, while the octupolar order in the negative domain is nearly unaffected.
-The slow evolution of $\langle\tau_{y}\rangle$ at late times, and the nearly unchanged $\langle\tau_{y}\rangle$ in the ‘--’ domain, indicate that heating effects are weak at these timescales.
-In order for this switching to occur, the phonon-induced field $(-\lambda^{2}Q_{0}^{2}/\Omega)\sin\phi$ must overcome the octupolar Weiss field in the ordered phase.
-We find that the phonon-induced field is too small to flip pseudospins deep inside a domain, but it is able to induce switching at domain walls where the Weiss field is strongly suppressed.
-Thus, negative domains must grow by flipping pseudospins at the boundary until the entire octupolar $\langle\tau_{y}\rangle$ flips.
-In Fig.3, we depict the time evolution of pseudospins under a two-phonon drive starting with an initial 3D domain wall configuration and a thermal population of phonon momenta.
-We focus on a single 2D slice through the lattice, with colours indicating positive $\langle\tau_{y}\rangle$ (red) and negative $\langle\tau_{y}\rangle$ (blue) and the black sites denoting the non-magnetic sublattice.
-One can clearly see the two-phonon drive inducing a shrinking of the $+\langle\tau_{y}\rangle$ domain as a function of time via spin flipping at the domain walls.
-In the bottom right panel of Fig.3 we show the expectation values of $\tau_{x},\tau_{y}$ and $\tau_{z}$ averaged over the lattice; we attribute the rapid oscillations to the high-frequency phonons and the apparent $x$-translational symmetry breaking to the initial distribution of phonon momenta.
-We provide similar pictures of the lattice spins for the switching presented in Fig.
-2(b) in the SM [72], as well as a simple theoretical model of this mechanism.
-Figure 3: Visualization of octupolar order switching starting with an initial domain wall configuration of pseudospins.
-We drive the $Q_{x}$ and $Q_{z}$ phonons with a relative phase of $\pi/2$ and show three time points around $t\sim\!7\hbar/J_{0}$, the point at which switching begins to occur.
-Here, we depict spins having $\langle\tau_{y}\rangle>0$ as red squares and $\langle\tau_{y}\rangle<0$ as blue squares.
-We initialize half the spins with $\langle\tau_{y}\rangle=+1.0$ and half with $\langle\tau_{y}\rangle=-1.0$.
-As time progresses, one can clearly see the shrinking of the positive (red) domain and growth of the negative (blue) domain as reflected in $\langle\tau_{y}\rangle$ averaged over the whole lattice, which is shown in the bottom right panel.
-Note that the rapid oscillation in the $\tau_{x},\tau_{y}$ and $\tau_{z}$ magnetization appears at the phonon frequency, while the apparent translational symmetry breaking is due to the initial phonon momenta, which are chosen from a thermal distribution.
-While experimental verification of our theoretical results on training and switching of octupolar order switching would necessitate a method of detecting the presence of octupolar order, this has proven notoriously difficult to do using conventional experimental techniques.
-Remarkably, Ref.[75] presents a promising route by which to unambiguously diagnose both the presence and, significantly, sign of the octupolar order parameter.
-Explicitly, rotational anisotropy second harmonic generation (SHG) measurements are conducted and the presence of octupolar order is characterized through interference between the time reversal even and odd channels, with the sign of the octupolar order determining the positions of the measured intensity maxima.
-Thus, we propose a tr-SHG experiment with a first measurement conducted at $t=0$ and a second conducted at time $t\sim 25\hbar/J_{0}$ ($\sim 16\rm ps$) looking for changes in the position of intensity maxima as a potential route to diagnose the training and switching of the octupolar order parameter.
-In the SM [72], we also provide further plots showing how our results depend on the phenomenological damping parameter $\eta$.
-Summary.— We have employed Monte Carlo and molecular dynamics simulations capable of modelling lattice degrees of freedom, in the form of local Einstein phonons, in addition to multipolar degrees of freedom to study Mott insulators which host non-Kramers doublets.
-Our equilibrium phase diagram for the model hosts ferro-octupolar or antiferro-quadrupolar ordered phases at low temperature, as partly described by previous work, but shows that the quadrupolar order is accompanied by lattice displacements.
-We next studied the out-of-equilibrium multipolar-phonon dynamics in the presence of a uniform excitation of Einstein phonon modes.
-Emulating pump-probe type experiments by giving a uniform impulsive kick to one phonon mode, we observed a characteristic response in the non-excited phonon below the octupolar ordering temperature which is absent in the disordered phase, leading to a dynamical signature of equilibrium octupolar order.
-Secondly, we model a continuous excitation of both phonons, with a relative phase between them, and observe the generation of octupolar order consistent with an effective field predicted by a lab frame Magnus expansion.
-This also led us to a proposed route to ultrafast picosecond timescale switching of octupolar order using phonons.
-Our work opens up unique directions in the study of Mott insulators playing host to multipolar moments, including potential routes by which to detect and control complex hidden orders.
+Using a drive which leads to a peak phonon amplitude $Q_{0}\!\sim\!1$-$3$ (i.e., $\sim\!1$-$3\%$ atomic displacement), we find that octupolar order in the positive domain is suppressed or can even switch sign, depending on the drive amplitude while the octupolar order in negative domains (not shown) is nearly unaffected.
+Furthermore, the slow evolution of $\langle\tau_{y}\rangle$ at late times indicate that heating effect of pseudospins is weak at these timescales, although additional cooling might be needed in experiments to extract the energy dissipated by the phonon modes.
+Nonlinear phononics experiments on CoF2 have realized large atomic displacements $\sim 2\%$ of Raman-active modes by driving an infrared-active phonon with $500$ fs pulse of 10MV/cm electric field [15]; such methods could potentially be extended to the osmates explored in this work.
+We find that octupolar switching is mediated by flipping of pseudospins at domain wall boundaries by the phonon-induced field $(-\lambda^{2}Q_{0}^{2}/\Omega)\sin\phi$; for $\phi\!=\!\pi/2$, this promotes growth of $\langle\tau_{y}\rangle\!<\!0$ domains.
+In Fig.3, we depict this time evolution of pseudospins under a two-phonon drive starting with an initial 3D domain wall configuration with periodic boundary conditions, and a thermal population of phonon momenta.
+We show a single 2D slice through the lattice, with colours indicating positive $\langle\tau_{y}\rangle$ (red) and negative $\langle\tau_{y}\rangle$ (blue) and the black sites denoting the non-magnetic sublattice.
+One can clearly see the two-phonon drive inducing a shrinking of the $+\langle\tau_{y}\rangle$ domain as a function of time.
+The bottom right panel of Fig.3 shows $\langle\tau_{y}\rangle$ averaged over the full lattice, evincing growth of octupolar magnetization; we attribute the superposed rapid oscillations to the high-frequency phonons.
+Additional data and an effective model of this mechanism is given in the SM [80].
+Figure 3: Visualization of octupolar order switching for an initial two-domain configuration (assuming periodic boundary conditions).
+We drive the $Q_{x}$ and $Q_{z}$ phonons with a relative phase $\phi\!=\!\pi/2$ for a short time $t=3\hbar/J_{0}$ ($\sim 1$ ps) and show configurations at representative times.
+Red and blue squares respectively depict pseudospins with $\langle\tau_{y}\rangle>0$ and $\langle\tau_{y}\rangle<0$.
+We clearly see the shrinking of the positive (red) domain and growth of the negative (blue) domain.
+This domain evolution is also reflected in the net octupolar magnetization $\langle\tau_{y}\rangle$ averaged over the whole lattice (bottom right panel) which is initially zero but evolves to negative values.
+Recent remarkable experiments [85] have shown that one can unambiguously diagnose the presence and, significantly, sign of ferro-octupolar order using rotational anisotropy second harmonic generation (SHG).
+The octupolar order is characterized through interference between the time reversal even and odd channels, with the sign of the octupolar order determining the positions of the measured intensity maxima [85].
+We propose a tr-SHG experiment which compares the position of intensity maxima of equilibrium versus the driven system at time $t\sim 3.0\hbar/J_{0}$ ($\sim 2.0\rm ps$) to diagnose the theoretically predicted build-up and switching of octupolar order.
+Additional plots showing dependence of our results on the phonon damping parameter $\eta$ are given in the SM [80].
+Summary.— We have developed computational tools to simulate the equilibrium phases and nonequilibrium dynamics of multipolar degrees of freedom coupled to dissipative Einstein phonons in multipolar Mott insulators.
+Our equilibrium phase diagram for the model hosts ferro-octupolar and quadrupolar orders at low temperature, as partly described by previous work, but additionally shows that the quadrupolar order is accompanied by weak lattice displacements.
+We have also shown how pump-probe techniques and two-mode phonon drive can uncover unique dynamical signatures and enable the ultrafast control of the ferro-octupolar state.
+This study opens up distinct directions in the detection and control of complex hidden orders in quantum materials.
 This research was funded by an NSERC Discovery Grant (AP), an Ontario Graduate Scholarship (KH), and an NSERC CGS-D fellowship (RS).
-We thank Sreekar Voleti for initial collaboration on this project.
-We acknowledge Dave Hsieh, Emily Zhang, William Bateman-Hemphill, and Cristian Batista for extremely helpful discussions.
+We thank Sreekar Voleti, David Hsieh, Prashant Padmanabhan, Emily Zhang, William Bateman-Hemphill, and Cristian Batista for extremely helpful discussions.
 Numerical computations were performed on the Niagara supercomputer at the SciNet HPC Consortium and the Digital Research Alliance of Canada.
+Data for the figures in this work is available, see [86].